Find the value of parameter for which the inequality 4x-a2x-a+3<=0 is satisfied by at least one real value of x.


Thanks!


 

Grade:11

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

Dear Palak J,

Ans:  your inequation is 4^x-a2^x-(a-3)<=0

make a hole square and we get,  [2^x-a/2]^2<=(a/2)^2  + a - 2

now the value of a square is >=0

hence the RHS term of the above equation is i.e (a/2)^2 + a - 2 >=0

solving we get,  a € R - (-6<=a<=2)

or a>=2 && a<=-6

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Palak J !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur